Dear Reader,
For The Past Week, My Math Teacher Assigns Us Algebra But I Don't Understand The Steps To Solving an Equation ex 5a-b=7. Help? :)

Found 2 solutions by Fombitz, Alan3354:
Answer by Fombitz(32388)

Which variable are you solving for a or b??

Answer by Alan3354(69443)

5a-b=7
----------
You have 2 variables, a and b. With only 1 eqn, you can solve for a in terms of b, and vice versa.
5a-b=7
Solve for a:
Add b to both sides
5a = b + 7
Divide by 5
a = (b+7)/5
-------------
Solve for b:
5a-b=7
Add b to both sides
5a = b+7
b+7 = 5a
Subtract 7
b = 5a - 7
----------
Solving an eqn is getting the specified variable on one side of the equal sign by itself.